### v1

# Why These Documents? Explainable Generative Retrieval with Hierarchical Category Paths

Abstract.
Generative retrieval has recently emerged as a new alternative of traditional information retrieval approaches. However, existing generative retrieval methods directly decode docid when a query is given, making it impossible to provide users with explanations as an answer for “Why this document is retrieved?”. To address this limitation, we propose Hierarchical Category Path-Enhanced Generative Retrieval (HyPE), which enhances explainability by generating hierarchical category paths step-by-step before decoding docid. HyPE leverages hierarchical category paths as explanation, progressing from broad to specific semantic categories. This approach enables diverse explanations for the same document depending on the query by using shared category paths between the query and the document, and provides reasonable explanation by reflecting the document’s semantic structure through a coarse-to-fine manner. HyPE constructs category paths with external high-quality semantic hierarchy, leverages LLM to select appropriate candidate paths for each document, and optimizes the generative retrieval model with path-augmented dataset. During inference, HyPE utilizes path-aware reranking strategy to aggregate diverse topic information, allowing the most relevant documents to be prioritized in the final ranked list of docids. Our extensive experiments demonstrate that HyPE not only offers a high level of explainability but also improves the retrieval performance in the document retrieval task.
Information Retrieval, Generative Retrieval, Explainability
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
1. Introduction
Figure 1. Existing generative retrieval methods fail to explain why specific documents are retrieved, as they directly decode docid (Upper). In contrast, our proposed HyPE method provides clear explanations by generating query-related hierarchical category paths leading to the docid (Lower).
Information retrieval (IR) systems play a vital role in helping users navigate the vast amount of online information. A fundamental task of these systems is document retrieval, which focuses on searching for and ranking documents that are relevant to a given query from a large corpus. Over the past few years, document retrieval has evolved rapidly and has become pivotal in several NLP downstream tasks, including web search engines, question answering, and dialogue systems. In the field of document retrieval, sparse retrieval (Robertson and Zaragoza, 2009; Formal et al., 2021) based on term-matching and dense retrieval (Karpukhin et al., 2020; Khattab and Zaharia, 2020; Xiong et al., 2021) based on semantic representations have been widely employed.
With the significant progress of language models, generative retrieval has recently emerged as a new paradigm in document retrieval. Generate retrieval aims to directly generate document identifier (i.e., docid) for a given query by leveraging pre-trained generative models such as BART (Lewis et al., 2020) and T5 (Raffel et al., 2020). This paradigm enables end-to-end optimization of the retrieval process, allowing for fine-grained interaction between the input query and docid, and significantly reduces memory usage by fully leveraging the parametric memory of a single pre-trained generative model.
Despite these advantages, generative retrieval continues to face the challenge of determining how to construct docid that effectively represent documents. As the docid serves as a representation of the entire document, defining one that accurately encapsulates the document’s contents is both crucial and challenging. Existing workds on generative retrieval have categorized docid into two types: semantic docid and lexical docid. Semantic docid represents each document as a unique identifier, either a single number or a series of numbers (e.g., ‘0-5-2’), derived from the dense representations encoded by a pre-trained text encoder (Devlin et al., 2019) and clustering techniques such as hierarchical k-means clustering (Tay et al., 2022; Wang et al., 2022; Sun et al., 2023) and product quantization (Zhou et al., 2022). On the other hand, lexical docid represents each document as human-readable text, such as titles (Cao et al., 2021), keywords (Zhang et al., 2023; Wang et al., 2023; Lee et al., 2023) and pseudo queries (Tang et al., 2023; Li et al., 2023).
Unlike semantic docid, lexical docid achieves interpretability by providing human-readable identifiers, but they still lack in explainability regarding why these documents were retrieved. In Figure 1 upper, existing generative retrieval methods directly decode docid when a query is given, making it impossible to provide users with explanations as an answer for “Why this document is retrieved?”. While users can recognize what a lexical docid represents, it is not always obvious why a certain docid is decoded for a specific query. The lack of explainability in retrieval systems can undermine the reliability of the retrieved documents and make it more challenging for users to explore additional information related to a specific query (Anand et al., 2022), making it a crucial issue to address. To address the aforementioned limitation, our research aims to design a generative retrieval framework that can provide retrieved results with reasonable explanations for the user’s query.
In this work, we propose Hierarchical Category Path-Enhanced Generative Retrieval (HyPE), which enhances explainability by generating hierarchical category paths step-by-step before decoding docid. Motivated by semantic docid, HyPE employs hierarchical category paths as explanation, progressing from broad to specific semantic categories. In Figure 1 lower, when queries about document “Dubai” are given, HyPE uses category paths like “Government ¿ Government by cities” or “Economy ¿ Economy by cities” to explain why document “Dubai” is retrieved for each query. This approach 1) enables diverse explanations for the same document depending on the query by using shared category paths between the query and the document, and 2) provides more reasonable and insightful explanation by reflecting the document’s semantic structure through a coarse-to-fine manner. Additionally, HyPE 3) can employ effective reranking of the retrieved results by leveraging multiple paths, which helps improve retrieval performance.
Specifically, HyPE consists of the following three steps: 1) constructing category paths based on an external high-quality semantic hierarchy and selecting appropriate candidate paths for each document using Large Language Models (LLM), 2) building a path-augmented dataset with candidate paths, and 3) optimizing a model with the path-augmented dataset. During inference phase, HyPE conducts a pseudo-reasoning [^1] Generating the hierarchical path resembles step-by-step reasoning. However, unlike natural language-based reasoning in LLM, we use the term “pseudo-reasoning” because the path structure is more akin to pseudo-code. process by generating the hierarchical category path step-by-step to decode docid, allowing it to serve as an explanation which enhances explainability. Additionally, HyPE employs path-aware reranking strategy, which simultaneously considers multiple pseudo-reasoning paths for each query. This strategy helps build a more robust retrieval system by capturing the semantic information of multiple category paths, thereby improving overall retrieval performance.
Our extensive experiments demonstrate that HyPE not only offers a high level of explainability but also improves the retrieval performance in document retrieval task. Additionally, HyPE can be applied orthogonally to various docid types (e.g., title, keywords), making it a versatile framework that can be seamlessly integrated into different generative retrieval systems. For reproducibility, our codes are publicly available at the anonymous github repository.222 https://github.com/augustinLib/HyPE-Why-These-Documents
The contributions of our work are summarized as follows:
  • We introduce HyPE, an explainable generative retrieval framework that generates query-specific hierarchical category paths for relevant documents before decoding their docid. These category path enables the retrieval system to provide users explanations of document retrieval.
  • We propose a new reranking strategy called path-aware reranking, which considers multiple category paths simultaneously to determine the final ranked list of docids.
  • We empirically show that HyPE improves both the explainability and accuracy of generative retrieval across various docid types, making it adaptable and easily integrable into different generative retrieval systems.
2. Related Work
2.1. Document Retrieval
Document retrieval is a representative task in information retrieval (IR) that involves searching for documents relevant to a given query within a document corpus. Traditionally, it has been accomplished through sparse retrieval and dense retrieval. Sparse retrieval (Robertson and Zaragoza, 2009; Formal et al., 2021) employs term-matching scoring methods, such as BM25 (Robertson and Zaragoza, 2009), to calculate relevance scores for retrieving documents. However, these methods struggle with the issue of lexical mismatch (Lin et al., 2020) due to terms such as homonyms and synonyms. To address these limitations, dense retrieval  (Karpukhin et al., 2020; Khattab and Zaharia, 2020; Xiong et al., 2021) has emerged with the recent development of pre-trained language models (PLMs). This approach utilizes encoder-based PLM, such as BERT (Devlin et al., 2019), to obtain dense representations for both queries and documents, and retrieve relevant documents through an Approximate Nearest Neighbor (ANN) index (Malkov and Yashunin, 2016). Although dense retrieval methods have demonstrated impressive performance by capturing deep semantic information that sparse retrieval methods cannot, it still has the limitation of not being able to optimize the model in an end-to-end manner.
2.2. Generative Retrieval
With the advancement of pre-trained language models, generative retrieval has emerged as an alternative to traditional document retrieval methods, establishing itself as a new paradigm. In contrast to traditional retrieval methods, generative retrieval leverages a single pre-trained generative model, such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020), to directly generate document identifier (docid) relevant to the query, enabling end-to-end optimization of the retrieval process. Additionally, this approach reduces reliance on external indexing, lowering the system’s demand for storage resources. Existing works have proposed various docid formats to represent the semantic information of documents, which can be categorized into two types: semantic docid and lexical docid. The semantic docid is represented as a sequence of numbers. These methods based on encoder-based PLM, like BERT (Devlin et al., 2019), to generate dense document embeddings. Afterward, techniques such as hierarchical k-means clustering (Tay et al., 2022; Wang et al., 2022; Sun et al., 2023) or product quantization (PQ) (Zhou et al., 2022) are applied to convert the document embeddings into sequential numeric form like “0-5-2”. This converted form serves as the semantic docid for the document, ensuring both efficient storage and effective retrieval based on the semantic structure, which is achieved by grouping semantically similar documents together. The lexical docid is a textual format that can well represent the semantic information of a document, and forms such as title (Cao et al., 2021), substrings (Bevilacqua et al., 2022), URL (Zhou et al., 2022), keywords (Wang et al., 2023; Zhang et al., 2023; Lee et al., 2023), and pseudo query (Tang et al., 2023; Li et al., 2023) are used to build the docid. Compared to methods that utilize semantic docid, methods based on lexical docid generate docid in a human-readable format, resulting in better interpretability. However, existing generative retrieval methods directly generate the docid for a user’s query, making it difficult to fully understand why the document is retrieved and how its rank is determined. To address these limitations, we propose a more reliable generative retrieval framework that not only generates relevant docid for the user’s query but also provides explanation.
Figure 2. Overview of HyPE framework. (1) HyPE constructs category paths using an external high-quality semantic hierarchy and employs LLM to select appropriate candidate paths for each document. (2) Then, HyPE links queries to the paths based on semantic relevance to construct path-augmented training set, and uses this to optimize the retrieval system. (3) During inference, HyPE employs path-aware reranking strategy to determine the final docid ranking by considering multiple paths.
3. Preliminaries
In this section, we formally define the task of generative retrieval and explain its overall process and relevant techniques.
3.1. Task Formulation
Given a corpus $\mathcal{C}=\{D_{1},D_{2},\dots,D_{n}\}$ where $D$ represents a document, generative retrieval aims to autoregressively generate the document identifier (i.e. docid) of the relevant document for a given query. To this end, the model is optimized for indexing task and retrieval task. The indexing task involves taking a document as the input and generating the corresponding docid, described by
(1) $\mathcal{M}^{\theta}(d\mid D)=\prod_{t=1}^{n}\mathcal{M}^{\theta}(d_{t}\mid D,d_{<t}),$
where $\mathcal{M}^{\theta}$ is a generative model, $D$ is a document, $d$ is the target docid, and $n$ is the token length of the target docid. The retrieval task focuses on processing a query as the input and generating the docid of a relevant document, described as follows:
(2) $\mathcal{M}^{\theta}(d\mid q)=\prod_{t=1}^{n}\mathcal{M}^{\theta}(d_{t}\mid q,d_{<t}),$
where $q$ is a query. In performing the aforementioned two tasks, it is important to 1) effectively represent the long document $D$ and 2) construct the docid $d$ that captures the overall semantic information of the document.
3.2. Document Representation and Identifier
3.2.1. Document representation
For the indexing task, each document is used as the input, making it crucial to define effective representations of the long document while preserving as much of its information as possible within the context length of the language model. The primary approaches to effectively representing documents are FirstP (Tay et al., 2022) and Document as Query (DaQ) (Wang et al., 2022). FirstP uses only the first $k$ tokens from the beginning of the document, while DaQ randomly extracts chunks from the entire document.
3.2.2. Document identifier
To ensure that docid effectively encodes semantic information of document, a variety of approaches have been proposed. As mentioned in section 2.2, docid can be broadly categorized into semantic docid and lexical docid. Semantic docid is a unique identifier, represented by either one number or a series of numbers. It is composed of document representations formed from encoder-based BERT, which are transformed into a compact numeric form. Lexical docid is a textual format designed to effectively convey the semantic content of a document. It can be constructed using various forms, such as the document’s title (Cao et al., 2021), substrings (Bevilacqua et al., 2022), keywords (Zhang et al., 2023; Lee et al., 2023; Wang et al., 2023), URL (Zhou et al., 2022), and pseudo query (Tang et al., 2023). Title and URL are used as docid directly from the dataset. Substrings are generated by the retrieval model using an FM index (Ferragina and Manzini, 2000), which creates specific n-grams within the document for retrieval. Keywords are extracted from the document using methods such as TF-IDF (Robertson and Walker, 1997), BM25 (Robertson and Zaragoza, 2009), or pre-trained language models (PLMs). Pseudo query is generated using query generation models such as docT5Query (Nogueira and Lin, 2020), and is then utilized as docid.
3.3. Optimization and Inference
3.3.1. Optimization via multi-task learning
Given a training dataset that consists of (query, document, docid), denoted by $\mathcal{X}=\{(q,D,d)\}$, the model is trained for both the indexing and retrieval tasks, maximizing the likelihoods in Equations (1) and (2), respectively:
(3) $\max_{\theta}\sum_{(q,D,d)\in\mathcal{X}}\mathcal{M}^{\theta}(d\mid D)+\sum_{(q,D,d)\in\mathcal{X}}\mathcal{M}^{\theta}(d\mid q).$
3.3.2. Indexing with synthetic query
In indexing task, documents are long and contain extensive information; on the contrary, in retrieval task, queries are relatively short and request specific information, highlighting a gap between the two tasks (Zhuang et al., 2023). To bridge this discrepancy, recent studies (Zhuang et al., 2023; Wang et al., 2022; Sun et al., 2023) have tried to integrate synthetic queries, generated by query generation (QG) models (Nogueira and Lin, 2020), into the training phase. The synthetic queries are helpful to improve the retrieval performance of generative retrieval models by effectively reducing the gap between queries and documents. Note that these synthetic queries are treated as alternative document representation, similar to FirstP and DaQ mentioned in 3.2.1, and are used as input for the indexing task (Pradeep et al., 2023; Zhuang et al., 2023; Sun et al., 2023).
3.3.3. Inference with constrained decoding
During inference, given an input query $q$, the model produces a top-$K$ ranked list of docids that have the largest likelihoods $\mathcal{M}^{\theta}(d\mid q)$. To ensure the generation of valid docids, the model employs constrained decoding, which mostly uses constrained beam search with a prefix trie (Cao et al., 2021).
4. Proposed Method
In this section, we present a novel framework, Hierarchical category Path-Enhanced generative retrieval (HyPE), which enhances explainability by generating hierarchical category paths step-by-step before decoding docid. Our hierarchical category paths serve as pseudo-reasoning process, providing a reasonable and reliable explanation for why this document is retrieved. Furthermore, HyPE employs path-aware reranking strategy, which simultaneously considers multiple pseudo-reasoning paths to determine the documents to retrieve. The overall framework is illustrated in Figure 2.
4.1. Candidate Path Set Construction
The first step of our HyPE framework is to construct a set of candidate hierarchical category paths for each document. In this context, the hierarchical category paths have to satisfy the following criteria to ensure explainability:
  • Semantic Hierarchy: Are they semantically hierarchical, allowing step-by-step progression in the generation process to clearly represent a specific semantic level?
  • Generalizability: Are they able to provide semantic information across a wide range of domains?
  • Specificity: Are they capable of sufficiently explaining specific and detailed semantic information?
To achieve this, we first construct the high-quality backbone hierarchy for category paths in open-domain. Then, for each document, we (1) filter out category paths based on semantic similarity calculated by a pre-trained text encoder, and (2) select several category paths that comprehensively represent the content of the document while specifically addressing certain topics within the document by the help of outstanding reasoning capabilities of LLM.
4.1.1. Hierarchical category path collection.
In the open-domain retrieval task, the category (or topic) hierarchy must encompass both a broad range of domain categories (i.e. width of tree) and sufficient semantic granularity (i.e. depth of tree) to ensure comprehensive and accurate retrieval system. To this end, we leverage Wikipedia’s category tree [^3] https://en.wikipedia.org/wiki/Special:CategoryTree as our backbone hierarchy of categories, setting the Main Topic classification category [^4] https://en.wikipedia.org/wiki/Category:Main_topic_classifications as the root node of the hierarchy. Considering the vast and complex nature of Wikipedia’s category tree, we limit the scraping process to a depth of four to construct our backbone hierarchy. Then, we linearize all the paths [^5] In this paper, for brevity, we use the term “path” to refer to a “category path”. within the hierarchy and convert them into a sequence of strings, thereby enabling more efficient processing and manipulation. The entire set of linearized category paths is denoted by $\mathcal{P}$. The statistics of collected hierarchical category paths are presented in Appendix A.1.
4.1.2. Candidate path set construction.
The next step is to utilize the knowledge of LLM to assign appropriate category paths to each document within the corpus. However, due to the context length of LLM, it is impossible to input all possible paths within the category hierarchy (collected in Section 4.1.1);
thus, we first filter out path set for each document $D$ by leveraging a bi-encoder. [^6] We use Sentence-T5 (Ni et al., 2022) as the bi-encoder and cosine similarity as similarity score. The pre-candidate path set is obtained by
(4) $\hat{\mathcal{P}}_{D}=\operatorname*{argTop-}_{p\in\mathcal{P}}k\text{ sim}(E(D),E(p)),$
where $E(\cdot)$ is the encoder, $\text{sim}(\cdot)$ is a similarity scoring function, and $k$ is the number of pre-candidate paths for each document. Then, we assign up to three paths that effectively represent the topics of the document by leveraging LLM few-shot prompting along with the document $D$ and its pre-candidate path set $\hat{\mathcal{P}}_{D}$  [^7] We use Llama-3-8B-Instruct as LLM..
(5) $\mathcal{P}_{D}=\text{LLM}({D},\hat{\mathcal{P}}_{D}).$
$\mathcal{P}_{D}$ refers to the final candidate path set of the document $D$. Our actual prompts used for this step can be found in Appendix A.4.
4.2. Optimization with Path-Augmented Dataset
The second step is to augment the existing training set $\mathcal{X}$ with path information, building a path-augmented training set $\mathcal{X}^{+}=\{(q,p^{q},D,d)\}$. To this end, we first (1) link each query to one of the document’s candidate paths based on semantic similarity computed by pre-trained text encoder, and then (2) use the query-path pairs along with the document-path pairs to optimize the generative retrieval model.
4.2.1. Linking Path with Query.
Using the candidate path set for each document, we build a path augmented training set $\mathcal{X}^{+}$. For each query-document pair in the training set $(q,D,d)\in\mathcal{X}$, we link the query $q$ to its most relevant path among the paths in the document’s candidate path set ${P}_{D}$. This query-path linking can be described as follows:
(6) $p^{q}=\operatorname*{argmax}_{p\in\mathcal{P}_{D}}\text{ sim}(E(q),E(p)),$
where $p^{q}$ is the path linked to the query $q$. This process is then applied to all queries in the training set. In the end, we construct the path-augmented training set, denoted by $\mathcal{X}^{+}=\{(q,p^{q},D,d)\}$.
4.2.2. Optimization.
By leveraging the path-augmented training set $\mathcal{X}^{+}$, we train our model $\mathcal{M}^{\theta}$ on both indexing and retrieval tasks, as described in 3.1. That is, the model is optimized by indexing loss and retrieval loss. First, the indexing loss is defined by
(7) $\mathcal{L}_{\text{indexing}}=-\sum_{p\in\mathcal{P}_{D}}\log\mathcal{M}^{\theta}(p,d\mid D)$
where $p$ is a path from $D$’s candidate path set. Note that there are multiple paths assigned to the document, so we optimize our model by using all candidate paths in $\mathcal{P}_{D}$. Through this process, the model can effectively memorize the content of the documents while mapping them to their relevant category paths. Next, the retrieval loss is described as follows:
(8) $\mathcal{L}_{\text{retrieval}}=-\log\mathcal{M}^{\theta}(p,d\mid q)$
The retrieval loss optimizes the model to learn the relationship between the query and its relevant path, allowing it to retrieve the documents that match the semantic of the query with the help of assigned paths. Finally, we define our total loss as the sum of the indexing loss and the retrieval loss, regarding the training process as a multi-task learning framework:
(9) $\mathcal{L}=\mathcal{L}_{\text{indexing}}+\mathcal{L}_{\text{retrieval}}.$
4.3. Inference with Path-Aware Reranking
During inference, HyPE generates the final ranked list of docids through two stages: 1) path generation stage and 2) docid decoding stage. First, in the path generation stage, our model $\mathcal{M}^{\theta}$ generates up to $K_{p}$ hierarchical category paths, each of which is denoted by $p_{j}$ for $j=1,\dots,K_{p}$, by using beam search; these are query-specific hierarchical category paths that encapsulate various topics related to the given query. Next, in the docid decoding stage, the model uses each generated hierarchical category path as the decoder’s input context and then applies constrained beam search to decode $m$ docids. For each path $p_{j}$, the model outputs $m$ number of docid-score pairs as follows:
(10) $Y_{j}=\{(d_{i},s_{i})\sim\ \mathcal{M}^{\theta}(\cdot\mid q,p_{j})\}_{i=1}^{m},$
where $s_{i}$ represents the score for the docid $d_{i}$ conditioned on the category path $p_{j}$. The remaining process is to aggregate $K_{p}$ number of docid-score pair sets for making the final ranked list of docids. At this point, we employ path-aware reranking strategy to remain only unique docid with the highest score, resulting in $\tilde{Y}$.
(11) $\tilde{Y}=\left\{(d,s)\mid s==\max\{s^{\prime}|(d,s^{\prime})\in Y_{j}\},\forall(d,s)\in\cup_{j=1}^{K_{p}}Y_{j}\right\}$
From the set of unique docid-score pairs, we obtain the final ranked list by sorting their scores in descending order, $Y_{\text{final}}=\text{sort}(\tilde{Y})$. By utilizing path-aware reranking strategy, HyPE can effectively capture the semantic information of an input query from multiple category paths, leading to improved retrieval performance.
5. Experiments
In this section, we design and conduct our experiments to answer the following research questions:
  • RQ1: Can HyPE effectively explain the semantic relationship between the query and document for document retrieval tasks?
  • RQ2: Can HyPE provide more effective and reasonable explanations compared to other methods for the user’s query?
  • RQ3: Can HyPE enhance the retrieval accuracy?
  • RQ4: Can path-aware reranking strategy of HyPE improve retrieval accuracy?
Table 1. Semantic relevance between query/explanation and document/explanation on 1,000 query-document pairs sampled from the NQ320K dev set, where each baseline successfully retrieves the ground truth document at rank 1. We also report BM25 as a sparse retrieval baseline.
Baseline Semantic Relevance
Query Document Overall
Title Docid 0.52 0.46 0.48
   + HyPE 0.50 0.56 0.52
Keyword Docid 0.42 0.54 0.47
   + HyPE 0.43 0.58 0.49
Summary Docid 0.46 0.69 0.55
   + HyPE 0.45 0.71 0.57
BM25 0.56 0.31 0.42
Figure 3. Examples of the document annotated for multiple queries in the NQ320K dev set. The generative retrieval model with HyPE generates query-specific category paths based on the topics of the document associated with each query, explaining why the document is retrieved for the particular query.
5.1. Experimental Settings
5.1.1. Dataset
We conduct our experiments on two public datasets, NQ320K (Kwiatkowski et al., 2019) and MS MARCO (Nguyen et al., 2016), which have been widely utilized in previous works (Tay et al., 2022; Wang et al., 2022). NQ320K is a dataset derived from Google’s Natural Questions (NQ) and comprises 307,373 query-document pairs along with 7,830 test queries, all linked to 109,739 documents sourced from Wikipedia. We divide the test set into two subsets, seen and unseen, following the setup in  (Wang et al., 2022; Sun et al., 2023), where the seen test includes queries whose annotated target documents are present in the training set, and the unseen test consists of queries with no labeled documents in the training set. On the other hand, MS MARCO comprises queries and web pages sourced from Bing search. We use the MS MARCO document ranking dataset, which contains 366,235 query-document relevant pairs and 5,187 test queries with 323,569 documents. The detailed statistics of the datasets are provided in Appendix 5.
5.1.2. Evaluation Metrics
We report Recall and Mean Reciprocal Rank (MRR) for NQ320K and MS MARCO. For NQ320K, we use Recall@{1, 10, 100} and MRR@100. For MS MARCO, we use Recall@{1, 10, 100} and MRR@10 as done in previous works (Sun et al., 2023; Wang et al., 2023).
Figure 4. Human evaluation of pairwise quality comparisons for retrieval explanations, generated by HyPE and baseline models (* : $p$-value ¡ 0.05).
Figure 5. An example from the NQ320K dev set showing that using only the title as the docid fails to retrieve the correct document. However, when HyPE is adopted, the generative retrieval model better understands the query and successfully retrieves the correct document.
Table 2. Retrieval accuracy of existing generative retrieval methods and our HyPE framework on the NQ320K dataset. * denotes the statistical significance on paired t-test $p<0.05$.
Method Full test Seen test Unseen test
R@1 R@10 R@100 M@100 R@1 R@10 R@100 M@100 R@1 R@10 R@100 M@100
GENRE (Cao et al., 2021) 55.2 67.3 75.4 59.9 69.5 83.7 90.4 75.0 6.0 10.4 23.4 7.8
[…]
+ HyPE 63.6 83.5 90.1 71.0 66.4 86.3 92.6 73.9 53.7 73.6 81.7 61.0
Improvement +2.3% +6.1% +2.5% +3.5% +2.5% +5.9% +2.8% +3.8% +1.1% +6.8% +1.6% +2.9%
NOVO (Wang et al., 2023) 69.3 89.7 95.9 76.7 - - - - - - - -
[…]
+ HyPE 60.7 79.1 86.2 67.6 66.6 84.6 90.7 73.4 40.1 60.2 70.6 47.5
Improvement -1.8% +2.6% +0.8% +0.0% -1.0% +2.8% +0.9% +0.5% -6.7% +2.0% +0.3% -2.7%
SEAL (Bevilacqua et al., 2022) 59.9 81.2 90.9 67.7 - - - - - - - -
[…]
+ HyPE 61.5 79.6 85.2 68.3 66.3 84.6 89.8 73.2 44.8 62.2 69.4 51.3
Improvement +1.0% +1.0% +1.3% +1.0% +0.9% +0.6% +1.4% +0.8% +1.8% +2.8% +1.3% +2.4%
GENRET (Sun et al., 2023) 68.1 88.8 95.2 75.9 70.2 90.3 96.0 77.7 62.5 83.6 92.5 70.4
[…]
+ HyPE 64.5 84.2 90.2 71.9 69.5 88.6 93.8 76.8 47.2 68.7 77.6 55.0
Improvement -1.2% +0.8% +1.0% -0.4% -1.0% +0.3% +0.3% -0.5% -2.9% +2.8% +3.6% +0.0%
5.1.3. Baselines
To validate the effectiveness of HyPE in generative retrieval, we conduct experiments on four representative docid types, introduced in Section 3.2.2, as our baseline generative retrieval approach.
  • Title docid (Cao et al., 2021) uses a document’s title as docid. For instances without a title, we use the first 16 tokens of the document as a pseudo-title, following the approach used in (Sun et al., 2023).
  • Keword docid (Wang et al., 2023) uses a sequence of keywords as docid that effectively represent the document. [^8] We use the off-the-shelf keyword extraction tool (Grootendorst, 2020). For NQ320K, we use 3 keywords, while for MS MARCO, we extract 5 keywords.
  • Summary docid uses the document summary as docid. [^9] We use the off-the-shelf text summarization model (Lewis et al., 2020). Although it has not been attempted before, a similar structure using substrings is employed in (Bevilacqua et al., 2022).
  • Atomic docid (Sun et al., 2023) uses a unique arbitrary integer as docid. We assign each document a unique arbitrary integer and generates a corresponding new token for it.
Existing generative methods employ various architectures and optimization techniques, which may introduce additional factors affecting performance. To isolate and specifically assess the impact of HyPE, we use the basic form of generative retrieval introduced in Section 3 as our main baseline. This approach allows us to focus on a direct comparison between plain docids and those enhanced with HyPE, ensuring that the observed effects are due to HyPE itself rather than other architectural or optimization variations. Furthermore, we additionally report the performance of existing methods to provide a reference for the performance of existing generative methods. We set a reference baseline for each docid type. Specifically, for each docid type, method that achieves the highest performance using the same dataset as ours is selected as the reference baseline for each docid. The detailed criteria for selecting the reference baseline can be found in 5.
5.1.4. Implementation Details
We use T5-base (Raffel et al., 2020) as our backbone model. For the input of the indexing task, we utilize the FirstP approach as our document representations (Section 3.2.1). Additionally, for the indexing task, we employ five synthetic queries, generated by using docT5query (Nogueira and Lin, 2020) with nucleus sampling with parameters $p=0.8$ and $t=0.8$. We use new [DOC] token to separate the path from the docid, which we insert between the path and the docid. We optimize our model as described in 4.2.2, while employing AdamW optimizer with a learning rate of 5e-4 and a batch size of 128, for up to 1M training steps. During the inference of HyPE, we adopt path-aware reranking strategy; for the path generation stage, we generates three category paths (i.e., $K_{p}=3$), and for the docid generation stage, we use constrained beam search with a beam size of 100 (i.e., $m=100$).
5.2. Explainability of HyPE (RQ1 & RQ2)
5.2.1. Quantitative analysis (RQ1).
We first quantitatively evaluate whether HyPE’s category path provides a valid explanation by effectively capturing the relationship between the query and the document. To do this, we use a model for the task of semantic textual similarity (STS), which measures the semantic relevance between two sentences (Agirre et al., 2012). Using the task, we assess the semantic relevance between the query and the explanation, as well as between the document and the explanation. Specifically, for each baseline, we use the model output as an explanation and calculate the STS scores for both the query-explanation and document-explanation pairs. We then compute the geometric mean of these two scores to evaluate how effectively the explanation captures the relationship between the query and the document. In this experiment, we also include BM25 as a baseline that can provide explanations for its retrieval results. For the explanation of BM25, we consider the top-3 terms with the highest BM25 scores between a query and a document.
As shown in Table 1, applying HyPE improves overall semantic relevance across all baselines. This indicates that HyPE’s category path effectively captures and explains the relationship between the query and the document. Notably, HyPE achieves higher overall relevance than the term-matching method (i.e., BM25), further proving the validity of the category path as an explanation.
Table 3. Retrieval accuracy of existing generative retrieval methods and HyPE framework on the MS MARCO dataset. * denotes the statistical significance on paired t-test $p<0.05$.
Method MS MARCO
R@1 R@10 R@100 M@10
NOVO (Wang et al., 2023) 49.1 80.8 92.5 59.2
[…]
+ HyPE 32.2 62.7 78.5 41.9
Improvement +1.6% +2.5% +1.7% +2.2%
SEAL (Bevilacqua et al., 2022) 25.9 68.6 87.9 40.2
[…]
+ HyPE 28.4 57.5 73.1 37.8
Improvement +1.1% +3.6% +2.2% +2.7%
GENRET (Sun et al., 2023) 47.9 80.8 91.6 59.2
[…]
+ HyPE 44.9 74.6 87.1 54.7
Improvement +2.3% +1.4% +1.8% +1.7%
5.2.2. Qualitative analysis (RQ2).
We evaluate the quality of explanations generated by HyPE through a human evaluation conducted via Amazon Mechanical Turk (AMT). We randomly sample 100 examples from the NQ320K dev set and ask three human judges to compare the quality of the explanations based on four distinct criteria: overall, specificity, reasonability and comprehensiveness. Detailed descriptions of these criteria are presented in Appendix A.3.
In Figure 4, HyPE outperforms both the title docid baseline and BM25 across all criteria, receiving high scores for its overall explanation of the retrieval process. Specifically, HyPE shows substantial margin of superiority in terms of specificity and reasonability. This demonstrates that HyPE provides clearer explanations of retrieval process, as well as more logical and reasonable explanation. Furthermore, HyPE beats other baselines in comprehensiveness, indicating that its hierarchical category path is effective in explaining not only narrow, specific details but also broader semantic information. These results highlight that HyPE’s pseudo-reasoning, which utilizes hierarchical category paths, provides users with a effective and reasonable explanation of the retrieval process.
Following human evaluation, we conduct two types of case studies to verify our HyPE provides more effective explanations for each query. Figure 3 illustrates HyPE’s pseudo-reasoning process in cases where a single document is annotated with multiple queries on different topics. In Example 1, for the query “the core of the sun in which the sun’s thermonuclear energy is produced”, the model generates paths related to the universe and energy conversion, clearly explaining the thematic relevance between the query and the document. However, for another query, “what stage of the star life cycle is the sun in”, it generates a path related to stellar evolution, which is different from the previously generated path but relevant to the query. This shows that HyPE can provide effective explanations to users by tailoring each query. Moreover, in Example 2, HyPE generates the same path for different queries with similar topics, specifically related to the building of the Statue of Liberty. This highlights HyPE ’s capability to provide coherent and consistent explanations for queries on related topics.
Figure 5 illustrates the difference between using only the title as the docid and considering paths alongside it with HyPE. For the query “Who sings love me tender in princess diaries 2?”, the baseline model using only the title fail to fully understand the query and retrieve an incorrect document [The Princess Diaries 2: Royal Engagement]. In contrast, when HyPE is adopted, the generative retrieval model first generates the path “entertainment ¿ theatre ¿ theatre soundtracks”, which captures the context related to “sing” that is missed by relying solely on the title. This allows the model to understand the query more specifically and successfully retrieve the correct document [The Princess Diaries soundtrack].
5.3. Retrieval Accuracy of HyPE (RQ3)
Table 2 shows retrieval accuracy of various docid types with HyPE on NQ320K, comparing the results against each baseline. Overall, HyPE consistently improves retrieval accuracy across all docid types in both seen test and unseen test. This demonstrates that HyPE’s hierarchical category paths can be orthogonally applied to enhance retrieval accuracy across different docid types, suggesting that integrating these paths into existing generative retrieval methods can further improve performance. While HyPE can be applied to all docid types effectively, the experimental results show that title docid yields the most significant performance improvement when HyPE is applied. Our paths, serve as a pseudo-reasoning allowing the model to navigate step-by-step through various semantic hierarchical categories before arriving docid. Since title docids are concise and inherently reflect a structured overview of a document, they aligns well with the HyPE’s hierarchical category paths, further enhancing retrieval accuracy and explanation clarity.
Figure 6. Performance changes of HyPE w.r.t. the number of decoded category paths to obtain a ranked docid list.
Additionally, we also report the experimental results on MS MARCO in Table 3. In contrast to NQ320K, MS MARCO, as a web search-based dataset, has varying document formats and contains significant semantic noise. Despite these challenges, the results show that HyPE consistently improves retrieval accuracy on this dataset as well. Although the hierarchical category paths are constructed using Wikipedia categories as the backbone, the consistent performance gains on MS MARCO emphasize the robustness and generalizability of HyPE. These findings suggest that HyPE can be widely applied to datasets across various domains in the future.
5.4. Analysis of Path-Aware Reranking (RQ4)
To validate the effectiveness of path-aware reranking strategy, we analyze the performance changes in retrieval accuracy with respect to the number of hierarchical category paths considered by HyPE. Figure 6 presents the analysis results, showing that retrieval accuracy improves as the number of paths increases across all docid types. Notably, there is a clear performance gap between the setting without path-aware reranking strategy (i.e., $K=1$) and with path-aware reranking strategy (i.e., $K>2$). These results strongly indicate that considering multiple paths through the path-aware reranking strategy allows the most relevant docids to be prioritized in the final ranked list, thereby enhancing retrieval accuracy. However, we observe that using too many paths eventually leads to a plateau in performance improvement. Beyond a certain threshold, additional paths tend to introduce noise or increase unnecessary computational complexity. Through our experiments, we can conclude that using three paths yields the most optimal retrieval accuracy for the majority of docid types.
6. Conclusion
This paper proposes HyPE, a novel framework designed to enhance the explainability of document retrieval by utilizing hierarchical category paths. By leveraging hierarchical category paths, HyPE addresses the limitations of existing generative retrieval methods that struggle with providing meaningful explanations for why specific documents are retrieved. Furthermore, we introduce path-aware reranking allows for the aggregation of diverse topic information, ensuring that the most relevant docids are prioritized in the final ranked list of docids. Our extensive experiments demonstrate that HyPE not only provides reliable and reasonable explanations but also enhances overall retrieval performance. To the best of our knowledge, this study is the first attempt to enhance explainability in generative retrieval, and we hope our research paves the way for positive advancements in this field.
Appendix A Appendix
A.1. Backbone category hierarchy
To address the criteria mentioned in Section 4.1—Semantic Hierarchy, Generalizability, and Specificity—we utilize Wikipedia’s category tree as the foundation for our hierarchical structure, designating the Main Topic classification category as the root node of the hierarchy. The Main Topic classification category is chosen as the root because it offers a broad and inclusive starting point, covering diverse domains. From this root node, our hierarchy branches into 40 child categories, each representing different fields such as Economy, Language, and Science, among others. These child categories, in turn, contain more specific subcategories that further detail their respective domains. To utilize Wikipedia’s category tree, we employed Selenium [^10] https://pypi.org/project/selenium/ to recursively scrape the Wikipedia and extract the Wikipedia category tree. When linearizing the category hierarchy into a hierarchical category path, each category is connected using the delimiter ¿. The delimiter ¿ is chosen among several candidate delimiters because it showed the highest semantic similarity to the natural language sentence “the right category is included in the left category”, as measured by Sentence-T5.
Level 1 Level 2 Level 3 Level 4 Total
40 1,330 13,383 95,240 109,993
Table 4. Statistics of the used category hierarchy, showing the number of nodes at each level (or depth).
A.2. Dataset overview
In this work, we use NQ320K and MS MARCO. For NQ320K, we follow NCI (Wang et al., 2022) setup and adhered to the seen and unseen test splits used in GENRET (Sun et al., 2023). For MS MARCO, we construct dataset based on the MSMARCO document ranking dataset, following setups from Ultron (Zhou et al., 2022), GENRET (Sun et al., 2023), and NOVO (Wang et al., 2023). Methods using the MSMARCO passage ranking dataset, such as TOME (Ren et al., 2023) and GLEN (Lee et al., 2023), or using NQ100K like SE-DSI (Tang et al., 2023), are excluded from consideration as reference baselines unless they report results specifically on the our dataset setup.
Dataset # Docs # Train queries # Test queries
NQ320K 109,739 307,373 7,830
MS MARCO 323,569 366,235 5,187
Table 5. Statistics of the document retrieval datasets used.
A.3. Human Evaluation
We assess the quality of the generated explanations by conducting a human evaluation, where we compare the outputs of HyPE to other baseline models using Amazon Mechanical Turk (AMT). We show the interface for the evaluation in Figure 7. We ask three human judges to evaluate the explanations based on the following four criteria.
  • Overall Which retrieval system output better explains the retrieval process overall?
  • Specificity Which retrieval system output provides more specific information?
  • Reasonability Which retrieval system output represents the retrieval process more logically and reasonably?
  • Comprehensiveness Which retrieval system output more comprehensively reflects the content of the document?
A.4. Prompts
Table 6 shows the prompt used to construct the path candidate set for the document with LLM.
Figure 7. Annotator interface of human evaluation on retrieval system output.
Table 6. The prompt for building final candidate path set.
Prompt: Select candidate path set for document
You’re a taxonomy expert.
You will receive a document along with a set of candidate taxonomy hierarchy paths for the document.
Your task is to select the path that can represent the document. Exclude paths that are too broad or less relevant or contain too specific information such as year.
You may list up to 3 paths, using only the paths in the candidate set. Do not include any explanation.
¡Document title¿
{Document title}
¡Document contents¿
{Document contents}
¡Candidate taxonomy hierarchy paths¿
{pre-candidate path set}
¡Selected representative taxonomy hierarchy paths¿
{Candidate path set}